# HD 44892: The youngest (or oldest?) gas-harbouring debris disc around an intermediate mass star

### v1

@@ 7 Summary and future prospects
Using ALMA, we have detected CO emission in a debris disc around a young, pre-MS intermediate-mass star HD 44892 with a fractional IR excess of 7.86 at 12 m (Iglesias et al., 2023), which falls within a range () where only one such object, HD 141569, has been previously identified with a measured CO gas mass. This provides a rare glimpse into the very stage of transition from protoplanetary to debris discs. HD 44892 is hence also one of the youngest known hybrid discs, along with the lower-mass (0.7 ) star NO Lup, with an age 1-3 Myr (Lovell et al., 2021a). For gas temperature in the range 20-50 K, the CO detection yields an estimate of CO gas mass of around 10-4 . On the other hand, our non-detections of 13CO and C18O lines provide upper limit on gas mass of around 10-2 , suggesting the CO gas mass we derive from 12CO emission should be treated as a lower limit.
Given the possibility of the primordial origin of gas, we also report the total gas mass values, calculated assuming an ISM ratio of H2= in the disc. This results in total H2 gas mass of 0.06 0.02 for 20 K gas and 0.12 0.01 for 50 K gas, yielding a gas-to-dust mass ratio of 3 at 20 K or 6 at 50 K. However, if this gas is of secondary rather than primordial origin, this ratio does not apply, since secondary gas is hydrogen-depleted (Smirnov-Pinchukov et al., 2022). In such a case, the gas is most likely CO dominated, making the CO gas mass a better representative of the total disc gas mass. Presence of circumstellar gas is further confirmed in our UVES observations of gas absorption, for which we were able to rule out interstellar origin.
@@
If this is correct, then the marginally resolved dust and gas extending at 40-45 au are very cold and vertically thin enough not to intercept stellar light - the only scenario to explain the lack of significant infrared emission from these spatial scales. Such flattened disc structures (and low-excess SEDs) may arise at late stages of protoplanetary disc evolution due to gas loss but also when all the very small, sub-micron dust has been removed (Panić & Min, 2017). Important to note is that removal of such dust creates conditions for gas survival beyond protoplanetary disc phase suggested by Nakatani et al. (2021). Assuming HD 44892 is an example of this scenario, future studies could involve scattered-light observations using telescopes such as the VLT/SPHERE, to verify if such small grains are indeed absent from its debris disc by, for example, if scattered light is absent, or has a colour indicative of scattering by larger grains (Mulders et al., 2013).
@@
The dust mass in the disc of HD 44892 is within an order of magnitude less than that of the hybrid disc HD 141569. The CO gas mass may be comparable to that of known gas-rich debris discs, depending on the optical thickness of 12CO emission. Given the very young age of HD 44892, primordial origin gas is likely (Nakatani et al., 2021; Smirnov-Pinchukov et al., 2022). Hence, it is possible that we have caught this disc just at the end of its primordial gas contents, which may explain the relatively low CO mass. This suggests HD 44892 is the next evolutionary stage following HD 141569, helping us understand the evolution of planetary systems better. Nonetheless, the young age of this object alone does not allow to make definitive conclusions about the origin of gas in this disc.
@@
It is important to note that from the CO gas mass alone, of (1.62 0.17) 10-4, it is impossible to fully rule out the secondary origin scenario. According to Cataldi et al. (2023), only CO-rich debris discs, which they classify as having M, cannot be solely explained by secondary models (like the one in Kral et al., 2017) and must still posses some primordial gas. On the other hand, they state that it is the debris discs with low CO gas masses, that are easily explainable by secondary means alone. Since the CO gas mass of HD 44892 is between the two, secondary or mixed origin is still possible. It is worth noting, however, that while models introduced in Kral et al. (2019) can explain the high CO gas masses in CO-rich debris discs, by introducing neutral carbon shielding and self-shielding of CO, more recent work of Marino et al. (2022) discusses the limitations of this scenario. In particular, they show that vertical mixing of gas in the debris disc reduces the effectiveness of shielding, limiting the accumulation of CO. Hence, for the most CO-rich systems secondary models alone are not yet able to explain this gas presence, and primordial origin remains the best available explanation. Cataldi et al. (2023) propose that observations of C i in debris discs where CO was found are needed to improve current models of gas, which might help answer questions regarding its origin. Since HD 44892 has a CO gas mass comparable to some of the debris discs with detected C i in their sample, it is possible that HD 44892 may also have detectable C i emission. However, given the low signal-to-noise ratio in our current observations, the likelihood of such a detection is low.
@@
If the alternative age of 800 Myr is considered, HD 44892 would be an outlier in terms of its dust mass (temperature-dependent), 12 m excess, and CO gas mas, strongly suggesting that this is the case of a young pre-main sequence star. The two proposed age scenarios, of 2.1 Myr and 800 Myr, correspond to different stellar mass estimates of 2.87 and 2.29 , respectively. Future, higher spatial and spectral resolution studies of CO emission with ALMA, could remove any remaining uncertainties, allowing for a dynamical mass estimation of the star.
@@ Acknowledgements
KS acknowledges funding from the Bell Burnell Graduate Scholarship Fund (BB0023). OP and DPI acknowledge support from the Science and Technology Facilities Council via grant number ST/X001016/1. TDP is supported by a UKRI Stephen Hawking Fellowship and a Warwick Prize Fellowship, the latter made possible by a generous philanthropic donation. JM acknowledges support from FONDECYT de Postdoctorado 2024 #3240612.
@@ Appendix A Debris discs with gas found through emission
V* CE Ant | 10 | 0.31[1] | M2Ve | 34.10 | 2.13 10-2 | 1.68 10-6 | Matrà et al. (2019)e | ||
HD 138813 | 10[3] | 2.15[3] | A0V | 136.60 | 1.30 10-1 | 1.92 10-5 | Lieman-Sifry et al. (2016)e | ||
HD 146897 | 10[3] | 1.32[3] | F2/3V | 132.19 | 2.14 10-1 | 2.90 10-6 | Lieman-Sifry et al. (2016)e | ||
@@
HD 121191 | 16[3] | 1.63[3] | A5IV/V | 132.29 | 9.02 10-3 | 1.99 10-4 | Moór et al. (2017)e | ||
HD 121617 | 16[3] | 1.90[3] | A1V | 117.89 | 1.30 10-1 | 3.86 10-2 | Moór et al. (2017)e | ||
HD 131488 | 16[4] | 1.8[6] | A1V | 152.24 | 3.21 10-1 | 4.75 10-2 | Moór et al. (2017)e | ||
HD 131835 | 16[3] | 1.81[3] | A2IV | 129.74 | 2.81 10-1 | 2.14 10-2 | Lieman-Sifry et al. (2016)e | ||
HD 156623 | 16[3] | 1.91[3] | A0V | 108.33 | 3.62 10-2 | 7.73 10-4 |
HD 181327 | 24 | 1.36 | F6V | 47.78 | 2.41 10-1 | 2.70 10-6 | Marino et al. (2016)e | ||
@@
HD 21997 | 42[3] | 1.83[3] | A3IV/V | 69.69 | 2.64 10-1 | 6.17 10-2 | Kóspál et al. (2013)e | ||
V* V1229 Tau | 170[5] | - | A0VpSi+Am | 138.31 | - | 6.33 10-4 | Péricaud et al. (2017)e | ||
HD 216956 | 440 | 1.9[6] | A4V | 7.70 | 1.49 10-2 | 1.89 10-7 | Matrà et al. (2017b)e | ||
* eta Crv | 1000-2000 | 1.39[3] | F2V | 18.24 | 1.07 10-2 | 5.92 10-7 | Marino et al. (2017)e |

### v2

@@ 7 Summary and future prospects
Using ALMA, we have detected CO emission in a debris disc around a young, pre-MS intermediate-mass star HD 44892 with a fractional IR excess of 7.86 at 12 m (Iglesias2023X-shooterEvolution), which falls within a range () where only one such object, HD 141569, has been previously identified with a measured CO gas mass. This provides a rare glimpse into the very stage of transition from protoplanetary to debris discs. HD 44892 is hence also one of the youngest known hybrid discs, along with the lower-mass (0.7 ) star NO Lup, with an age 1-3 Myr (Lovell2021ALMADispersal). For gas temperature in the range 20-50 K, the CO detection yields an estimate of CO gas mass of around 10-4 . On the other hand, the non-detections of 13CO and C18O lines provide upper limits on gas mass of around 10-2 . Since we do not know whether the 12CO emission we detected is optically thin or thick, the CO gas mass we derived from 12CO emission should be treated as a lower limit, and may be up to 10-2 , which is the upper limit derived from 13CO non-detection.
Given the possibility of the primordial origin of gas, we also report the total gas mass values, calculated assuming an ISM ratio of H2= in the disc. This results in total H2 gas mass of 0.06 0.02 for 20 K gas and 0.12 0.01 for 50 K gas, yielding a gas-to-dust mass ratio of 3 at 20 K or 6 at 50 K. The H2 masses presented are based on 12CO emission, which may be optically thick, and hence represent lower limits to the total disc gas mass. However, if this gas is of secondary rather than primordial origin, this ratio does not apply, since secondary gas is hydrogen-depleted (Smirnov-Pinchukov2022LackGas). In such a case, the gas is most likely CO dominated, making the CO gas mass a better representative of the total disc gas mass.
Presence of circumstellar gas is further confirmed in our UVES observations of gas absorption, for which we were able to rule out interstellar origin. From the UVES data we also derived an accretion rate for HD 44892 of = -7.64 and = -7.69 for the two observed epochs. This value is comparable to that of HD 141569 exceeding the level expected for CO-rich debris discs and is similar to values found in protoplanetary discs. If the gas was secondary, then based on the M⊕ CO gas mass upper limit, from our 13CO non-detection, it would be depleted on very short timescales, and replenishment at the required rate seems unlikely. This strongly suggests that at least the inner disc retains primordial, -rich gas.
@@
If this is correct, then the marginally resolved dust and gas extending at 40-45 au are very cold and vertically thin enough not to intercept stellar light - the only scenario to explain the lack of significant infrared emission from these spatial scales. Such flattened disc structures (and low-excess SEDs) may arise at late stages of protoplanetary disc evolution due to gas loss but also when all the very small, sub-micron dust has been removed (Panic2017EffectsLocation). Important to note is that removal of such dust creates conditions for gas survival beyond protoplanetary disc phase suggested by Nakatani2021PhotoevaporationRemnants. Assuming HD 44892 is an example of this scenario, future studies could involve scattered-light observations using telescopes such as the VLT/SPHERE, to verify if such small grains are indeed absent from its debris disc by, for example, if scattered light is absent, or has a colour indicative of scattering by larger grains (Mulders2013WhyHD100546).
@@
The dust mass in the disc of HD 44892 is within an order of magnitude less than that of the hybrid disc HD 141569. The CO gas mass may be comparable to that of known gas-rich debris discs, depending on the optical thickness of 12CO emission. Given the very young age of HD 44892, primordial origin gas is likely (Nakatani2021PhotoevaporationRemnants; Smirnov-Pinchukov2022LackGas). Hence, it is possible that we have caught this disc just at the end of its primordial gas contents, which may explain the relatively low CO mass. This suggests HD 44892 is the next evolutionary stage following HD 141569, helping us understand the evolution of planetary systems better. Nonetheless, the young age of this object alone does not allow to make definitive conclusions about the origin of gas in this disc.
@@
It is important to note that from the CO gas mass alone, of (1.62 0.17) 10-4, it is impossible to fully rule out the secondary origin scenario. According to Cataldi2023PrimordialALMA, only CO-rich debris discs, which they classify as having M, cannot be solely explained by the current secondary models (like the one in Kral2017PredictionsPlanetesimals). A solution may be either an improvement of secondary models of gas production or presence of primordial gas. On the other hand, they state that it is the debris discs with low CO gas masses, that are easily explainable by secondary means alone. Since the CO gas mass of HD 44892 is between the two, secondary or mixed origin is still possible. It is worth noting, however, that while models introduced in Kral2019ImagingDiscs can explain the high CO gas masses in CO-rich debris discs, by introducing neutral carbon shielding and self-shielding of CO, more recent work of Marino2022VerticalLifetime discusses the limitations of this scenario. In particular, they show that vertical mixing of gas in the debris disc reduces the effectiveness of shielding, limiting the accumulation of CO. Hence, for the most CO-rich systems secondary models alone are not yet able to explain this gas presence, and primordial origin remains the best available explanation (Cataldi2023PrimordialALMA; Marino2022VerticalLifetime). Cataldi2023PrimordialALMA propose that observations of C i in debris discs where CO was found are needed to improve current models of gas, which might help answer questions regarding its origin. Since HD 44892 has a CO gas mass comparable to some of the debris discs with detected C i in their sample, it is possible that HD 44892 may also have detectable C i emission in higher sensitivity observations.
@@
If the alternative age of 800 Myr is considered, HD 44892 would be an outlier in terms of its dust mass (temperature-dependent), 12 m excess, and CO gas mass, strongly suggesting that this is the case of a young pre-main sequence star. The two proposed age scenarios, of 2.1 Myr and 800 Myr, correspond to different stellar mass estimates of 2.87 and 2.29 , respectively. Future, higher spatial and spectral resolution studies of CO emission with ALMA, could remove any remaining uncertainties, allowing for a dynamical mass estimation of the star.
@@ Acknowledgements
KS acknowledges funding from the Bell Burnell Graduate Scholarship Fund (BB0023). OP and DPI acknowledge support from the Science and Technology Facilities Council via grant number ST/X001016/1. TDP is supported by a UKRI Stephen Hawking Fellowship and a Warwick Prize Fellowship, the latter made possible by a generous philanthropic donation. JM acknowledges support from FONDECYT de Postdoctorado 2024 #3240612 and ANID – Millennium Science Initiative Program – Center Code NCN2024_001.
@@
We thank the anonymous referee for their insightful comments and suggestions, which provided valuable perspectives on the nature of this system.
@@ Appendix A Debris discs with gas found through emission
| V* CE Ant | 10 | 0.31[1] | M2Ve | 34.10 | 2.13 10-2 | 1.68 10-6 | Matra2019On7e | ||
| HD 138813 | 10[3] | 2.15[3] | A0V | 136.60 | 1.30 10-1 | 1.23 10-3 | Lieman-Sifry2016DebrisALMAe | ||
| HD 146897 | 10[3] | 1.32[3] | F2/3V | 132.19 | 2.14 10-1 | 2.90 10-6 | Lieman-Sifry2016DebrisALMAe | ||
@@
| HD 121191 | 16[3] | 1.63[3] | A5IV/V | 132.29 | 9.02 10-3 | 5.83 10-3 | Moor2017MolecularStarse | ||
| HD 121617 | 16[3] | 1.90[3] | A1V | 117.89 | 1.30 10-1 | 3.86 10-2 | Moor2017MolecularStarse | ||
| HD 131488 | 16[4] | 1.8[6] | A1V | 152.24 | 3.21 10-1 | 4.75 10-2 | Moor2017MolecularStarse | ||
| HD 131835 | 16[3] | 1.81[3] | A2IV | 129.74 | 2.81 10-1 | 2.14 10-2 | Lieman-Sifry2016DebrisALMAe | ||
| HD 156623 | 16[3] | 1.91[3] | A0V | 108.33 | 3.62 10-2 | 6.48 10-4 |
| HD 181327 | 24 | 1.36 | F6V | 47.78 | 2.41 10-1 | 2.70 10-6 | Marino2016ExocometaryRinge | ||
@@
| HD 21997 | 42[3] | 1.83[3] | A3IV/V | 69.69 | 2.64 10-1 | 6.17 10-2 | Kospal2013ALMA21997e | ||
| V* V1229 Tau | 170[5] | - | A0VpSi+Am | 138.31 | - | 6.33 10-4 | Pericaud2017TheDiskse | ||
| HD 216956 | 440 | 1.9[6] | A4V | 7.70 | 1.49 10-2 | 1.89 10-7 | Matra2017DetectionCometse | ||
| * eta Crv | 1000-2000 | 1.39[3] | F2V | 18.24 | 1.07 10-2 | 5.92 10-7 | Marino2017ALMAPlanetse |